# Understanding the Potential of FPGA-Based Spatial Acceleration for Large Language Model Inference

### 23 Dec 2023 v1

@@ 4.3. Multi-Device Performance Estimation
Figure 9. Latency estimations of Vicuna-13B model on multiple FPGAs.
@@ 5.1. HLS Kernels
Linear Operators. Linear operators are the key operation in Transformer models because they are ubiquitous and compute-intensive. There are two types of linear operations in the Transformer layers, activation-weight matrix multiply (A-W GEMM) with bias, and activation-activation (A-A GEMM) matrix multiply. A-W GEMM includes the projection layers and the linear layers in the FFN, with weights buffered on-chip and activations streamed from an input FIFO; A-A GEMM includes the two GEMM operators in the SDP attention, with one of the input activations stored in a double buffer and the other one streamed.
@@
Each MAC unit can be implemented with a single DSP block and can provide one-MAC-per-cycle throughput. Based on the discussion in §4.2.2, we adopt the W4A8 quantization scheme for our accelerator design, which maximizes the utilization of available resources. As a result, the matrix multiplications involve either int4 by int8 or int8 by int8 operations, which are too large for LUTs, thus relying primarily on DSPs or specialized compute blocks (e.g., AIE (Xilinx, 2022a)). In AMD FPGAs, the DSP48E2 hard blocks can support 18-bit by 27-bit multiplication and accumulation (Xilinx, 2021), allowing to pack two multiplications into one slice for a W4A8 quantized model to save DSP resources and achieve a larger $M$. Figure 10(b) shows the bit packing method for 4-bit by 8-bit integer multiplications. One activation is filled into the lower 8 bits of the 18-bit DSP input, and two weights are filled into 0-to-3 and 13-to-16 bit positions of the 27-bit DSP input to avoid overlapping results. Finally, the two multiplication results are extracted by bit-slicing the 45-bit DSP result. Notice that since the DSP output is wide enough, we can also pack two 8-bit by 8-bit integer multiplications into one DSP slice by further offsetting the second weight and output. With DSP packing, we can easily double $M$ to achieve higher performance but with much fewer DSPs.
Non-Linear Operators. Since quantizing non-linear operations can lead to a significant degradation in model accuracy (Shen et al., 2020; Xiao et al., 2023), and these non-linear operations are not the bottleneck of the design, we directly implement the floating-point version of the operators in HLS. Specifically, we buffer a row of elements for softmax and LayerNorm functions, which requires conducting reduction along the last dimension. Consequently, this approach eliminates the need to wait for the complete results for the computation of these non-linear operators and effectively prevents dataflow stalling.
Figure 11. Overall dataflow architecture of a single Transformer layer that uses post-LayerNorm scheme (Radford et al., 2019). SDP denotes scaled dot-product. Orange nodes denote the GEMM kernels. Yellow nodes are the non-linear kernels, including softmax (SM), LayerNorm (LN), and GELU (GL). Green rectangles represent the FIFOs between kernels, and purple rectangles are the data loaders.
@@ 5.2. Accelerator Design
We integrate the proposed kernels in §5.1 to create a high-performance hardware accelerator. The overall architecture of our proposed accelerator is depicted in Figure 11. Different operations including linear operations and non-linear operations are connected with FIFOs, except that the KV cache discussed in §3.2.2 leverages double buffers. This design only reads the input tensors from off-chip memory, and stores the results back to memory at the end of each layer. The intermediate activation values are directly streamed to the next operation. $L_{\cdot}$ are data loaders that are responsible for loading data from memory or streaming buffers. Given the contemporary trend of multi-die FPGA designs (Guo et al., 2021; Du et al., 2022), explicit dataflow partitioning becomes necessary to meet timing requirements. Our target device, Alveo U280 FPGA (Xilinx, 2021), has three chip dies called super logic regions (SLRs). Consequently, we partition the dataflow into three regions to confine each region fully inside SLR.
We leverage the proposed analytical framework to guide our accelerator design. Since typical Transformer models have $d_{\text{FFN}}=4d$ (Devlin et al., 2018; Radford et al., 2019) and $l<d$, according to work balancing of Equation (11), we have $M_{q,k,v,p}=M$, $M_{a_{1},a_{2}}<M$, and $M_{f1,f2}=4M$. A straight-forward division is to put the PEs for $q$, $k$, $v$, SDP, and $p$ on SLR0, $f1$ on SLR1, and $f2$ on SLR2 so that each SLR roughly contains $4M$ MAC units. However, we observe that scaling up the linear operators in FFN poses significant challenges to timing closure. Among various configurations of systolic arrays we tested, the maximum capacity of one SLR at 250 MHz is three of 8 $\times$ 16 systolic arrays; a single 16 $\times$ 16 one fails timing. Therefore, we only leverage 8 $\times$ 8 and 8 $\times$ 16 systolic arrays for simplicity. To minimize the number of SLR crossings, we put $q$, $k$, and $v$ on SLR0 and use 8 $\times$ 8 systolic arrays. MHA and the $p$ projection are on SLR1, with $a_{1}$ and $a_{2}$ using 8 $\times$ 8, and $p$ using 8 $\times$ 16 systolic arrays. $f_{1}$ and $f_{2}$ operators on SLR2 using 8 $\times$ 16 systolic arrays. Therefore, it can still form a relatively balanced 3:4:4 resource utilization ratio.
@@ 6. Evaluation on FPGAs
In this section, we implement two design points studied in §4 to validate the feasibility of our framework. We first describe our experimental setup and perform a single FPGA evaluation. Later, we project the results to multiple FPGAs.
@@ 6.1. Experiment Setup
We test the publicly available BERT and GPT2 models listed in Table 2. We conduct post-training quantization and use the W4A8 quantization scheme for BERT and the W8A8 scheme for GPT, which are prevalent settings in nowadays LLM inference (Zhao et al., 2023; Xiao et al., 2023). Unless specially specified, we set the sequence length $l$ to $128$.
We implement and run the actual design on an Alveo U280 FPGA (Xilinx, 2021) with $M=256$ for the kernels. This FPGA has 4032 BRAM18K blocks, 9024 DSP slices, 2.6M flip-flops, 1.3M LUTs, and 960 URAM blocks. It has three SLRs with almost equal amount of resources. All the kernels are implemented in C++ with Vitis HLS v2022.1 (Xilinx, 2022f), and synthesized with Vivado backend toolchains. We use OpenCL with Xilinx RunTime (XRT) for hardware execution and Xilinx Board Utility (xbutil) for computing power measurements. The environment for GPU experiments is listed in §4.1, and NVIDIA system management interface (nvidia-smi) is used for measuring GPU power. Notice the quantized models on GPUs are slower than the FP16 models, as the quantization methods normally leverage fake quantization and lack high-performance GPU kernels to support efficient inference. Therefore, we directly compare our accelerator with the best FP16 GPU results. The FPGA on-board results match the outputs from the quantized model in PyTorch and are able to achieve the same accuracy. The latency results are the average across fifty runs.
@@ 6.2. On-Board Evaluation
Table 4. Experimental results compared with other FPGA-based accelerators. FTrans uses the RoBERTa model but the model architecture and configuration are the same as the BERT-base model in Table 2. Sequence lengths are set as 512.
@@
FQ-BERT (Liu et al., 2021) ZCU111 214 W4A8 95.16 10.51 3.66$\times$ 679 (31%) 3287 (77%) 201K (24%) 190K (45%) N/A
FTrans (Li et al., 2020a) VCU118 N/A Fixed16 419.15 2.39 16.11$\times$ N/A 6531 (96%) 507K (N/A) 451K (17%) N/A
@@
We first compare our BERT accelerator with FQ-BERT (Liu et al., 2021) and FTrans (Li et al., 2020a), two FPGA-based accelerators for BERT-variant models. To make a fair comparison, we employ the same W4A8 quantization precision with FQ-BERT and assess the model accuracy using the CoLA dataset (Devlin et al., 2018), a widely-used language understanding task. The fp16 model achieves an accuracy of 56.84%, while our W4A8 quantized model attains 56.76% accuracy. We only compare the performance of the core encoder layer, and the latency results are directly obtained from the original papers. From Table 4, we can see our spatial architecture delivers a much lower latency and a higher throughput with a much lower DSP usage. Even though our evaluation device is not exactly the same as the baselines, our throughput improvement still surpasses the resource increment of FF and LUT. Specifically, our accelerator is 3.66$\times$ faster than FQ-BERT (Liu et al., 2021) and 16.11$\times$ faster than FTrans (Li et al., 2020a). This is mainly due to the overlapped computation of different stages, memory and DSP packing, and higher frequency because of careful floorplanning. Moreover, our analytical framework precisely predicts the performance of the accelerator with less than 2ms of differences, showing the practicality of our approach.
We next design an accelerator for the GPT2 model. We support importing quantized models from different quantization frameworks (Xiao et al., 2023; Zhao et al., 2023; Chee et al., 2023). Specifically, we export the W8A8 model from SmoothQuant (Xiao et al., 2023) and achieve 62.2% on the LAMBADA dataset (Paperno et al., 2016), whereas the FP16 model demonstrates an accuracy of 65.0%. We compare our GPT accelerator with the state-of-the-art GPT accelerator, DFX (Hong et al., 2022), which employs a temporal architecture with an instruction set and uses the same U280 FPGA device for on-board evaluation. On average, we are 2.16$\times$ and 1.10$\times$ faster than DFX in the prefill and decode stage respectively. This is because our spatial architecture overlaps the computation and greatly eliminates off-chip memory access. We can also see our estimations in §3 align closely with the actual performance, achieving a 92% accuracy for the prefill stage. For the decode stage, the estimated latencies are lower than the actual results, which is mainly because the initial interval between two operators is not significantly smaller than the execution time of one stage, contributing to a notable increase in latency.
@@ 6.3. Ablation Study
We further investigate the efficiency of our kernel functions. We conduct experiments on our template systolic array function with the AutoSA-generated systolic array, which is a highly optimized state-of-the-art systolic array implementation (Wang et al., 2021). From Table 5, we can see our implementation achieves the same level of performance compared to AutoSA. While maintaining the same DSP usage, the resource usage of our kernel function is much smaller than AutoSA. This also explains the reason why our analytical model can precisely predict the performance of spatial accelerators. The presence of latency-predictable kernels as foundational components plays a pivotal role in this predictability. Additionally, our function offers enhanced customizability, accommodating varying sizes and the choice of different quantization schemes.
@@
Table 5. Performance of our systolic array library function. The GEMM kernel is extracted from the first FFN layer in the BERT-base model with size $(512,768)\times(768,3072)$. We use a 16 $\times$ 16 systolic array to calculate the int8 GEMM. The theoretical peak performance without DSP packing is $(512\times 768\times 3072)/(16\times 16)$ cycles ×$\times 3.33$ ns/cycle = $15.71$ ms.
  Latency (ms) BRAM DSP FF LUT
  Ours (w/o DSP packing) 15.73 0 (0%) 256 (2%) 88284 (3%) 168190 (12%)
Ours (w/ DSP packing) 15.73 0 (0%) 128 (1%) 79969 (3%) 244439 (18%)
AutoSA (Wang et al., 2021) 15.71 514 (12%) 256 (2%) 100138 (3%) 244032 (18%)

### 07 Apr 2024 v2

@@ 4.3. Multi-Device Performance Estimation
Figure 9. Latency estimations of Vicuna-13B model on multiple FPGAs.
@@ 5.1. HLS Kernels
Linear Operators. Linear operators are the key operators in Transformer models because they are ubiquitous and compute-intensive. There are two types of linear operators in the Transformer layers, activation-weight matrix multiply (A-W GEMM) with bias, and activation-activation (A-A GEMM) matrix multiply. A-W GEMM includes the projection layers and the linear layers in the FFN, with weights buffered on-chip and activations streamed from an input FIFO; A-A GEMM includes the two GEMM operators in the SDP attention, with one of the input activations stored in a double buffer and the other one streamed.
@@
Each MAC unit can be implemented with a single DSP block and can provide one-MAC-per-cycle throughput. Based on the discussion in §4.2.2, we adopt the W4A8 quantization scheme for our accelerator design, which maximizes the utilization of available resources. As a result, the matrix multiplications involve either int4 by int8 or int8 by int8 operations, which are too large for LUTs, thus relying primarily on DSPs or specialized compute blocks (e.g., AIE (xilinxAIE, )). In AMD FPGAs, the DSP48E2 hard blocks can support 18-bit by 27-bit multiplication and accumulation (u280, ), enabling the packing of two multiplications into one slice for a W4A8 quantized model to save DSP resources and achieve a larger $M$. Figure 10(b) shows the bit packing method for 4-bit by 8-bit integer multiplications. One activation is filled into the lower 8 bits of the 18-bit DSP input, and two weights are filled into 0-to-3 and 13-to-16 bit positions of the 27-bit DSP input to avoid overlapping results. Finally, the two multiplication results are extracted by bit-slicing the 45-bit DSP result. Notice that since the DSP output is wide enough, we can also pack two 8-bit by 8-bit integer multiplications into one DSP slice by further offsetting the second weight and output. With DSP packing, we can easily double $M$ to achieve higher performance but with much fewer DSPs.
Non-Linear Operators. Since quantizing non-linear operators can lead to a significant degradation in model accuracy (sheng2020qbert, ; xiao2023smoothquant, ), and these non-linear operators are not the bottleneck of the design, we directly implement the floating-point version of the operators in HLS. Specifically, we buffer a row of elements for softmax and LayerNorm functions, which requires conducting reduction along the last dimension. Consequently, this approach eliminates the need to wait for the complete results for the computation of these non-linear operators and effectively prevents dataflow stalling.
Figure 11. Overall dataflow architecture of a single Transformer layer that uses post-LayerNorm scheme (radford2019gpt2, ). SDP denotes scaled dot-product. Orange nodes denote the GEMM kernels. Yellow nodes are the non-linear kernels, including softmax (SM), LayerNorm (LN), and GELU (GL). Green rectangles represent the FIFOs between kernels, and purple rectangles are the data loaders.
@@ 5.2. Accelerator Design
We integrate the proposed kernels in §5.1 to create a high-performance hardware accelerator. The overall architecture of our proposed accelerator is depicted in Figure 11. Different operators including linear and non-linear operators are connected with FIFOs, except that the KV cache discussed in §3.2.2 leverages double buffers. This design reads input tensors from off-chip memory and stores the results back to memory after each layer. Intermediate activation values are directly streamed to the next operator. Initially, all parameters are stored in DRAM, and data loaders $L_{(\cdot)}$ are responsible for loading data from DRAM or streaming buffers. Storing parameters off-chip reduces the number of FPGA devices needed and avoids the need to build different bitstreams for different network layers. After completing a layer, the accelerator fetches new parameters from the host to the device for the subsequent layer. Since data loading is hidden from the computation, overall latency remains unaffected. Given the contemporary trend of multi-die FPGA designs (guo2021autobridge, ; du2022hisparse, ), explicit dataflow partitioning becomes necessary to meet timing requirements. Our target device, Alveo U280 FPGA (u280, ), has three chip dies called super logic regions (SLRs). Consequently, we partition the dataflow into three regions to confine each region fully inside SLR. According to our placement constraints, AMD Vitis toolchain will automatically insert AXI Register Slice IPs to pipeline SLR crossing.
We leverage the proposed analytical framework to guide our accelerator design. Since typical Transformer models have $d_{\text{FFN}}=4d$ (devlin2018bert, ; radford2019gpt2, ) and $l<d$, according to work balancing of Equation (11), we have $M_{q,k,v,p}=M$, $M_{a_{1},a_{2}}<M$, and $M_{f1,f2}=4M$. A straight-forward division is to put the PEs for $q$, $k$, $v$, SDP, and $p$ on SLR0, $f1$ on SLR1, and $f2$ on SLR2 so that each SLR roughly contains $4M$ MAC units. However, we observe that scaling up the linear operators in FFN poses significant challenges to timing closure. Among various configurations of systolic arrays we tested, the maximum capacity of one SLR at 250 MHz is three of 8 $\times$ 16 systolic arrays; a single 16 $\times$ 16 one fails timing. Therefore, we only leverage 8 $\times$ 8 and 8 $\times$ 16 systolic arrays for simplicity. We also explore using LUT-based multipliers as they provide greater flexibility for placement compared to DSPs. However, the presence of additional inter-LUT wires results in a much lower frequency (191 MHz) compared to the DSP-based multipliers. To minimize the number of SLR crossings, we put $q$, $k$, and $v$ on SLR0 and use 8 $\times$ 16 systolic arrays, which also ensures a relatively low latency for the first stage based on Equation (8). MHA and the $p$ projection are on SLR1, with $a_{1}$ and $a_{2}$ using 8 $\times$ 8, and $p$ using 8 $\times$ 16 systolic arrays. $f_{1}$ and $f_{2}$ operators on SLR2 using 8 $\times$ 16 systolic arrays. Therefore, it can still form a relatively balanced 3:2:2 resource utilization ratio for linear operators.
@@ 6. Evaluation on FPGAs
In this section, we implement two design points studied in §4 to validate the feasibility of our framework. We first describe our experimental setup and perform evaluation on a single FPGA.
@@ 6.1. Experiment Setup
We test the publicly available BERT and GPT2 models listed in Table 2. We conduct post-training quantization and use the W4A8 quantization scheme for BERT and the W8A8 scheme for GPT, which are prevalent settings in nowadays LLM inference (zhao2023atom, ; xiao2023smoothquant, ).
Figure 12. Physical layout of the implemented spatial accelerator.
We implement and run the actual design on an Alveo U280 FPGA (u280, ) with $M=256$ for the kernels. This FPGA has 4032 BRAM18K blocks, 9024 DSP slices, 2.6M flip-flops, 1.3M LUTs, and 960 URAM blocks. It has three SLRs with an almost equal amount of resources. All the kernels are implemented in C++ with Vitis HLS v2022.1 (vitis_hls_2022, ), and synthesized with Vivado backend toolchains. Figure 12 shows the final device layout of the implemented accelerator. We use OpenCL with Xilinx RunTime (XRT) for hardware execution and Xilinx Board Utility (xbutil) for computing power measurements. The environment for GPU experiments is listed in §4.1, and NVIDIA system management interface (nvidia-smi) is used for measuring GPU power. Notice the quantized models on GPUs are slower than the FP16 models, as the quantization methods normally leverage fake quantization and lack high-performance GPU kernels to support efficient inference. Therefore, we directly compare our accelerator with the best FP16 GPU results. The FPGA on-board results match the outputs from the quantized model in PyTorch and are able to achieve the same accuracy. The latency results are the average across fifty runs.
@@ 6.2. On-Board Evaluation
Table 4. Experimental results compared with other FPGA-based accelerators. Sequence lengths are set as 512.
@@
- SLR0 - - - 4.86 - - 130 (19%) 482 (17%) 200K (23%) 167K (38%) 3 (1%)
- SLR1 - - - 14.63 - - 136 (20%) 590 (19%) 240K (28%) 212K (49%) 50 (16%)
- SLR2 - - - 19.81 - - 123 (18%) 708 (23%) 213K (25%) 191K (44%) 58 (18%)
FQ-BERT (liu2021fqbert, ) ZCU111 214 W4A8 95.16 10.51 3.66$\times$ 679 (31%) 3287 (77%) 201K (24%) 190K (45%) N/A
TRAC (patrick2022trac, ) ZCU106 200 Fixed8 347.48 2.88 13.36$\times$ 181 (29%) 1379 (80%) 128K (28%) 126K (55%) N/A
@@
We first compare our BERT accelerator with FQ-BERT (liu2021fqbert, ) and TRAC (patrick2022trac, ), two FPGA-based accelerators for the BERT model. To make a fair comparison, we employ the same W4A8 quantization precision with FQ-BERT and assess the model accuracy using the CoLA dataset (devlin2018bert, ), a widely-used language understanding task. The fp16 model achieves an accuracy of 56.84%, while our W4A8 quantized model attains 56.76% accuracy. We only compare the performance of the core encoder layers, and report the best on-board latency results of the baselines obtained from the original papers. Both baselines use a sequence length of 128, so we scale their latency results to match our sequence length of 512. FTrans (li2020ftrans, ) also targets BERT-variant models, but it does not provide frequency and sequence length in the paper, so we cannot make a direct comparison. From Table 4, we can see our spatial architecture delivers a much lower latency and a higher throughput with a much lower DSP usage. Even though our evaluation device is not exactly the same as the baselines, our throughput improvement still surpasses the resource increment of FF and LUT. Specifically, our accelerator is 3.66$\times$ faster than FQ-BERT (liu2021fqbert, ) and 13.36$\times$ faster than TRAC (patrick2022trac, ). Compared to temporal architectures, our spatial architecture can efficiently overlap the execution of the operators in the model and eliminate most of the data movement overhead. In our design, one layer can start operation once the previous layer finishes computing one tile of the feature map, which typically takes only tens of cycles. In contrast, temporal architectures need to wait for the entire tensor to be produced, which may take hundreds of cycles. Therefore, even if the per-layer latency of spatial architectures is longer, the end-to-end latency can still be significantly lower than the temporal architectures employed by FQ-BERT and TRAC. Furthermore, our analytical framework precisely predicts the performance of the accelerator with less than 2ms of differences, showing the practicality of our approach.
We next design an accelerator for the GPT2 model. We support importing quantized models from different quantization frameworks (xiao2023smoothquant, ; zhao2023atom, ; chee2023quip, ). Specifically, we export the W8A8 model from SmoothQuant (xiao2023smoothquant, ) and achieve 62.2% on the LAMBADA dataset (denis2016lambada, ), whereas the FP16 model demonstrates an accuracy of 65.0%. We compare our GPT accelerator with the state-of-the-art GPT accelerator, DFX (hong2022dfx, ), which employs a temporal architecture with an instruction set and uses the same U280 FPGA device for on-board evaluation. On average, we are 2.16$\times$ and 1.10$\times$ faster than DFX in the prefill and decode stage respectively. This is because our spatial architecture overlaps the computation and greatly eliminates off-chip memory access. We can also see our estimations in §3 align closely with the actual performance, achieving a 92% accuracy for the prefill stage. For the decode stage, the estimated latencies are lower than the actual results, which is mainly because the initial interval between two operators is not significantly smaller than the execution time of one stage, contributing to a notable increase in latency.
@@ 6.3. Ablation Study
We begin by examining the latency of different SLRs. As shown in rows 3-5 in Table 4, the overall latency of the BERT accelerator is nearly the sum of the latency of SLR0 and SLR2, aligning with the pipeline diagram in Figure 4. Moreover, the computation in SLR1 largely overlaps with that of SLR2 due to the fully pipelined design. The resource usage across different SLRs is also similar, resulting in a balanced design.
We further investigate the efficiency of our kernel functions. We conduct experiments on our template systolic array function with the AutoSA-generated systolic array, which is a highly optimized state-of-the-art systolic array implementation (wang2021autosa, ). From Table 5, we can see our implementation achieves the same level of performance compared to AutoSA. While maintaining the same DSP usage, the resource usage of our kernel function is much smaller than AutoSA. Since the prediction of a single GEMM kernel is accurate and can achieve the theoretical performance, our analytical model can precisely predict the performance of spatial accelerators when combining multiple linear operators. The presence of latency-predictable kernels as foundational components plays a pivotal role in this predictability. Additionally, our function offers enhanced customizability, accommodating varying sizes and the choice of different quantization schemes.
@@
Table 5. Latency and resource usage of our systolic array library function. Results are directly derived from the HLS report in 300MHz. The GEMM kernel is extracted from the first FFN layer in the BERT-base model with size $(512,768)\times(768,3072)$. We use a 16 $\times$ 16 systolic array to calculate the int8 GEMM. The theoretical peak performance without DSP packing is $(512\times 768\times 3072)/(16\times 16)$ cycles ×$\times 3.33$ ns/cycle = $15.71$ ms.
  Latency (ms) Effective $M$ BRAM DSP FF LUT
  Ours (w/o DSP packing) 15.73 256 0 (0%) 256 (2%) 88284 (3%) 168190 (12%)
Ours (w/ DSP packing) 15.73 128 0 (0%) 128 (1%) 79969 (3%) 244439 (18%)
AutoSA (wang2021autosa, ) 15.71 N/A 514 (12%) 256 (2%) 100138 (3%) 244032 (18%)
Lastly, we analyze the performance of the non-linear operators. As shown in Table 6, we observe that the softmax operator in the MHA module incurs the highest latency, primarily due to the need to compute the exponential function. Since these operators are elementwise and only require a row of data to start the computation, they can be easily fused with the preceding linear operators in the pipeline, thereby not significantly impacting the overall latency. For instance, the combined latency of a GEMM kernel (10.77ms) and the softmax operator (6.67ms) greatly exceeds the latency of SLR1 in Table 4 (14.63ms///300MHz×$\times$245MHz), indicating substantial overlap between the softmax operator and other operators. Again, these ablation studies show that considering only the linear operators in the analytical framework is sufficient to achieve an accurate latency estimation.
Table 6. Performance and resource usage of non-linear operators in our kernel library. Kernel sizes are set to match those of the BERT model in Table 4. Results are directly derived from the HLS report in 300MHz.
  Operator Latency (ms) BRAM DSP FF LUT
  Softmax 6.67 8 38 4835 7447
LayerNorm 0.85 20 80 18751 12746
GeLU 0.67 0 256 26193 16472